Evaluate the area between y = x2 and the line y = 2x.


1
Expert's answer
2021-04-14T10:57:00-0400

To find the area under the curve "y=x^2 \text{ and } y=2x" , we sketch the curve.




to get the upper limit and lower limit "[a,b]" of the functions, we get the point of intersection.


Since "y=x^2 \text{ and } y=2x" , then


"x^2=2x\\
\implies x^2-2x=0\\
x(x-2)=0\\
x=0 \text{ or } x-2=0\\
\therefore x= 0,2"

We proceed to integrate the function about the limits so as to get the area under the curve:


"A= \int_{0}^{2} (2x-x^2)\,dx\\
= \Bigg [x^2-\frac{x^3}{3} \Bigg ]^2_1\\
\qquad \qquad \quad= \Bigg [2^2-\frac{2^3}{3} \Bigg ]-\Bigg [0^2-\frac{0^3}{3} \Bigg ]\\
= 4-\frac{8}{3}\\
A = 2.67 \text{ sq.units}"